# Bridging Neural ODE and ResNet: A Formal Error Bound for Safety Verification

### v1

@@ Abstract
Keywords:
Neural ODE, ResNet, Formal relationship, Safety verification, Reachability analysis
@@ 1 Introduction
Neural ordinary differential equations (neural ODE) are gaining prominence in continuous-time modeling, offering distinct advantages over traditional neural networks, such as memory efficiency, continuous-time modeling, adaptive computation balancing speed and accuracy, as well as enhanced interpretability by modeling dynamics via differential equations [5, 12, 21]. This surge in interest stems from recent advancements in differential programming, which have enhanced the ability to model complex dynamics with greater flexibility and precision [22].
@@
Figure 1: Illustration of the proposed framework to verify Model $1$ based on the outcome of the verification of Model $2$ and a bound $\varepsilon$ on the maximal error between the models.
@@
Abstraction-based verification (i.e., verifying properties of one model by working on an abstraction of its behaviors into a simpler model) has been a popular topic in the past decades outside of the AI field [24]. On the other hand, it has rarely been used up to now within the field of AI verification. One of its main current applications is on the topic of neural network model reduction, where the verification of a neural network is achieved at a lower computational cost on a reduced network with less neurons, see e.g. [4] for unidirectional relationships, or [26] for bidirectional ones through the use of approximate bisimulation relations. Although the overall principle of the proposed approach in our paper is similar (abstracting a model by one that over-approximates the set of all its behaviors), the main difference with the above works between two discrete neural networks is that our paper considers the formal relationships between a continuous neural ODE model and a discrete ResNet one.
@@ 2.1 Neural ODE and ResNet models
with state $x\in\mathbb{R}^{n}$, initial state $x(0)=u$, and vector field $f:\mathbb{R}^{n}\rightarrow\mathbb{R}^{n}$ defined as a finite sequence of classical neural network layers (such as fully connected layers, convolutional layers, activation functions, batch normalization). The state trajectories of (1) are defined as:
@@ 3.1 Lagrange remainder
Proposition 1 (Lagrange remainder)
@@ 3.2 Error function
In our context of Section 2, the function $f$ is assumed not to be explicitly dependent on the depth $t$ due to its definition as a single residual block with classical layers. Therefore, the partial derivative $\frac{\partial f(x(t))}{\partial t}$ is equal to 00, and the third term of (5) thus reduces to:
@@ 3.3 Bounding the error set
$\mathcal{R}_{\varepsilon}(\mathcal{X}_{in})$ =$=\left\{\Phi(1,u)-(u+f(u))~{}|~{}u\in\mathcal{X}_{in}\right\}$
=$=\left\{\left.\frac{1}{2}f^{\prime}(\Phi(t^{*},u))f(\Phi(t^{*},u))~{}\right|~{}t^{*}\in[0,1],~{}u\in\mathcal{X}_{in}\right\}.$ (9)
@@
The error set can then be re-written based on this reachable tube, as follows:
$\mathcal{R}_{\varepsilon}(\mathcal{X}_{in})$ =$=\left\{\left.\frac{1}{2}f^{\prime}(x)f(x)~{}\right|~{}x\in\mathcal{R}^{\text{tube}}_{\text{neural ODE}}(\mathcal{X}_{in})\right\}$
@@
The next step, in line 2 of Algorithm 1, is to over-approximate this error set $\mathcal{R}_{\varepsilon}(\mathcal{X}_{in})$. One possible approach to achieve this is to define the static function $\varepsilon=\frac{1}{2}f^{\prime}(x)f(x)$ and apply to it some set-propagation techniques (such as interval arithmetic [11], Taylor models [17], or affine arithmetic [6]) to bound the set of output errors $\varepsilon$ corresponding to any state $x\in\Omega^{\text{tube}}_{\text{neural ODE}}(\mathcal{X}_{in})$ in the reachable tube over-approximation. An alternative approach, which provided a tighter error bounding set in the particular case of the numerical example presented in Section 4, is to re-define this static error function as a discrete-time nonlinear system $x^{+}=\frac{1}{2}f^{\prime}(x)f(x)$, and then use existing reachability analysis toolboxes to over-approximate the reachable set of this system after a single time step. Note that in this case, it is important that this final reachable set is computed as a single step, and not decomposed into a sequence of smaller intermediate steps whose iterative updates of the internal state would have no mathematical meaning for the static (stateless) error function.
@@
where $L$ is a Lipschitz constant of the neural ODE vector field. The term $\left\|\frac{1}{2}f^{\prime}(x)f(x)\right\|_{\infty}$ can be obtained by first over-approximating the error set by $\Omega_{\varepsilon}(\mathcal{X}_{in})$ in the same way we did, but the infinity norm forces to expand this set to make it symmetrical around 00, and then keeping only the maximum value among its components (thus corresponding to a second expansion of this set into an hypercube whose width along all dimensions is the largest width of the previous set). In addition, for any system with non-zero Lipschitz constant, the factor $\frac{e^{L}-1}{L}$ is always greater than $1$, which increases this error bound even more.
@@ 3.4 Verification proxy
We first focus on the case of Algorithm 1 to verify the safety property on the neural ODE, based on the reachability analysis of the ResNet. This first verification proxy relies on the set-based version of (6):
@@
stating that the reachable output set of the neural ODE is included in the output set over-approximation of the ResNet $\Omega_{\text{ResNet}}(\mathcal{X}_{in})$, expanded by the bounding set of the error $\Omega_{\varepsilon}(\mathcal{X}_{in})$ obtained after applying the first two lines of Algorithm 1 as described in Section 3.3.
Therefore, this verification procedure is achieved as in Algorithm 1, by first using existing set-propagation or reachability analysis tools to compute an over-approximation ${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\Omega_{\text{ResNet}}(\mathcal{X}_{in})}$ of the ResNet output set (line 3). Then in line 4, an over-approximation of the neural ODE output set can be deduced from (11) by taking the set addition of ${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\Omega_{\text{ResNet}}(\mathcal{X}_{in})}$ and our error bound ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\Omega_{\varepsilon}(\mathcal{X}_{in})}$. If $\Omega_{\text{neural ODE}}(\mathcal{X}_{in})$ is contained in the safe set ${\color[rgb]{0,0.6,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.6,0}\pgfsys@color@cmyk@stroke{0.64}{0}{0.95}{0.40}\pgfsys@color@cmyk@fill{0.64}{0}{0.95}{0.40}\mathcal{X}_{s}}$, then the neural ODE satisfies the safety property, otherwise the result is inconclusive (line 5-9).
@@
2:  compute the over-approximation of the error set ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\Omega_{\varepsilon}(\mathcal{X}_{in})}$, $\forall x\in\Omega^{\text{tube}}_{\text{neural ODE}}(\mathcal{X}_{in})$;
3:  compute the over-approximation of the ResNet output ${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\Omega_{\text{ResNet}}(\mathcal{X}_{in})}$;
@@
5:  if $\Omega_{\text{neural ODE}}(\mathcal{X}_{in})\subseteq{\color[rgb]{0,0.6,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.6,0}\pgfsys@color@cmyk@stroke{0.64}{0}{0.95}{0.40}\pgfsys@color@cmyk@fill{0.64}{0}{0.95}{0.40}\mathcal{X}_{s}}$ then
@@
2:  compute the over-approximation of the negative error set ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\Omega_{-\varepsilon}(\mathcal{X}_{in})}$, $\forall x\in\Omega^{\text{tube}}_{\text{neural ODE}}(\mathcal{X}_{in})$;
@@
5:  if ${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\Omega_{\text{ResNet}}(\mathcal{X}_{in})}\subseteq{\color[rgb]{0,0.6,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.6,0}\pgfsys@color@cmyk@stroke{0.64}{0}{0.95}{0.40}\pgfsys@color@cmyk@fill{0.64}{0}{0.95}{0.40}\mathcal{X}_{s}}$ then
@@ 4.2 Computing the error bound
Using CORA [1], we compute the error bound $\Omega_{\varepsilon}(\mathcal{X}_{in})$ from Theorem 3.1 as follows. First, we over-approximate the reachable tube of the neural ODE $\mathcal{R}^{\text{tube}}_{\text{neural ODE}}$ over the time interval $[0,1]$ as a sequence of zonotopes, where each zonotope corresponds to an intermediate time range. For each zonotope in the reachable tube, we apply discrete-time reachability analysis on the error function (7) at $t=1$, treating it as a discrete-time nonlinear system in a single step without intermediate time splitting. This results in a new zonotope that over-approximates the error set starting from that particular reachable tube zonotope. The total error set is thus guaranteed to be contained in the union of these error zonotopes across all time steps. To simplify its use in the safety verification experiments in Section 4.3, we compute the interval hull of this union, yielding a hyperrectangle that over-approximates ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\Omega_{\varepsilon}}(\mathcal{X}_{in})$ illustrated in Figure 2 in red, and showing 20 error zonotopes in different colors, corresponding to the error bound of each intermediate time range used in the reachable tube.
@@
To contextualize our proposed error bound, we compare it with the error bound proposed in [23]. For that, we first compute the infinity norm of our error set $\|\Omega_{\varepsilon}(\mathcal{X}_{in})\|_{\infty}=0.064$, which corresponds to a positive and scalar bound on the error, thus implying that its set representation in the state space (represented in yellow in Figure 3) is necessarily symmetrical around 00 and with the width that is identical on all dimensions (since the infinity norm takes the largest width across all dimensions). The set-based error bound ($\Omega_{\varepsilon}(\mathcal{X}_{in})$ represented in red) obtained from our method is thus always contained in this infinity norm.
Next, we compute the Lipschitz constant for the vector field of the neural ODE $L=\|\tau+W\|_{\infty}=3.62$, and then we obtain the error bound in [23] as $\frac{(e^{L}-1)}{L}\|\Omega_{\varepsilon}(\mathcal{X}_{in})\|_{\infty}=0.64$. This final error bound, represented in magenta in Figure 3, is 10 times wider (on each dimension) than the infinity norm of our error set in yellow, and about $16$ millions times larger (in volume over the $5$-dimensional state space) than our error set $\Omega_{\varepsilon}(\mathcal{X}_{in})$ in red. The improved tightness of our proposed approach is therefore very significant.
@@ 4.3 Experiments on safety verification
Using the error bound computed in Section 4.2, we can verify safety properties for the neural ODE output set based on the ResNet output set and the error bound set (i.e., $\Omega_{\text{ResNet}}(\mathcal{X}_{in})+\Omega_{\varepsilon}(\mathcal{X}_{in})$), or vice versa for the ResNet output set based on the neural ODE output set and the negative error bound set (i.e., $\Omega_{\text{neural ODE}}(\mathcal{X}_{in})+\Omega_{-\varepsilon}(\mathcal{X}_{in})$).
In Figure 4, we compute the over-approximation of the ResNet output set ${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\Omega_{\text{ResNet}}}$ using simple bound propagation through the ResNet function with CORA. By adding the error bound $\Omega_{\varepsilon}$, we obtain a zonotope (shown in red) that is guaranteed to contain $\mathcal{R}_{\text{neural ODE}}(\mathcal{X}_{in})$. The figure also includes black points representing neural ODE outputs for random initial conditions in $\mathcal{X}_{in}$, with their convex hull (black set) approximating the true reachable set $\mathcal{R}_{\text{neural ODE}}(\mathcal{X}_{in})$. Since the safe set ${\color[rgb]{0,0.6,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.6,0}\pgfsys@color@cmyk@stroke{0.64}{0}{0.95}{0.40}\pgfsys@color@cmyk@fill{0.64}{0}{0.95}{0.40}\mathcal{X}_{s}}$ contains the over-approximation ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\Omega_{\text{ResNet}}(\mathcal{X}_{in})+\Omega_{\varepsilon}(\mathcal{X}_{in})}$, we guarantee that the neural ODE true reachable set is safe, as:
${\color[rgb]{0,0.6,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.6,0}\pgfsys@color@cmyk@stroke{0.64}{0}{0.95}{0.40}\pgfsys@color@cmyk@fill{0.64}{0}{0.95}{0.40}\mathcal{X}_{s}}\supseteq{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\Omega_{\text{ResNet}}(\mathcal{X}_{in})+\Omega_{\varepsilon}(\mathcal{X}_{in})}\supseteq\mathcal{R}_{\text{neural ODE}}.$
From Figure 4, we can see that the ResNet and neural ODE reachable sets are very similar due to the ResNet role as a discretization of the neural ODE, but they are not identical. Indeed, some neural ODE outputs (black points) lie outside ${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\Omega_{\text{ResNet}}}$, highlighting the necessity of the error bound $\Omega_{\varepsilon}(\mathcal{X}_{in})$ to ensure that the over-approximation captures all possible neural ODE outputs.
@@
Conversely, in Figure 5, we compute the over-approximation of the neural ODE reachable set ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\Omega_{\text{neural ODE}}(\mathcal{X}_{in})}$. By adding the negative error bound $\Omega_{-\varepsilon}$, we obtain a zonotope (shown in red) that encapsulates ${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathcal{R}_{\text{ResNet}}(\mathcal{X}_{in})}$. Similarly, the figure includes blue points representing ResNet outputs for random inputs in $\mathcal{X}_{in}$, with their convex hull (blue set) approximating the true reachable set ${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathcal{R}_{\text{ResNet}}(\mathcal{X}_{in})}$. Since the safe set ${\color[rgb]{0,0.6,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.6,0}\pgfsys@color@cmyk@stroke{0.64}{0}{0.95}{0.40}\pgfsys@color@cmyk@fill{0.64}{0}{0.95}{0.40}\mathcal{X}_{s}}$ is a super set that contains the over-approximation ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\Omega_{\text{neural ODE}}(\mathcal{X}_{in})+\Omega_{-\varepsilon}(\mathcal{X}_{in})}$, we guarantee that the ResNet true reachable set is safe, as:
${\color[rgb]{0,0.6,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.6,0}\pgfsys@color@cmyk@stroke{0.64}{0}{0.95}{0.40}\pgfsys@color@cmyk@fill{0.64}{0}{0.95}{0.40}\mathcal{X}_{s}}\supseteq{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\Omega_{\text{neural ODE}}(\mathcal{X}_{in})+\Omega_{-\varepsilon}(\mathcal{X}_{in})}\supseteq{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathcal{R}_{\text{ResNet}}}.$

### v2

@@ 1 Introduction
Neural ordinary differential equations (neural ODE) are gaining prominence in continuous-time modeling, offering distinct advantages over traditional neural networks, such as memory efficiency, continuous-time modeling, adaptive computation balancing speed and accuracy [5, 14, 23]. This surge in interest stems from recent advancements in differential programming, which have enhanced the ability to model complex dynamics with greater flexibility and precision [24].
@@
Figure 1: Illustration of the proposed framework to verify Model 11 based on the outcome of the verification of Model 22 and a bound $\varepsilon$ on the maximal error between the models.
@@
Abstraction-based verification (i.e., verifying properties of one model by working on an abstraction of its behaviors into a simpler model) has been a popular topic in the past decades outside of the AI field [27]. Within the field of AI verification, its primary application has been on abstracting specific model components rather than the whole model itself, as in approaches based on convex relaxation of nonlinear ReLU activation functions [13, 11]. On the other hand, full-model abstraction has been mostly unexplored for AI verification, except on the topic of neural network model reduction, where the verification of a neural network is achieved at a lower computational cost on a reduced network with less neurons, see e.g. [4] for unidirectional relationships, or [29] for bidirectional ones through the use of approximate bisimulation relations. Although the overall principle of the proposed approach in our paper is similar (abstracting a model by one that over-approximates the set of all its behaviors), the main difference with the above works between two discrete neural networks is that our paper considers the formal relationships between a continuous neural ODE model and a discrete ResNet one.
@@ 2.1 Neural ODE and ResNet models
with state $x\in\mathbb{R}^{n}$, initial state $x(0)=u$, and vector field $f:\mathbb{R}^{n}\rightarrow\mathbb{R}^{n}$ defined as a finite sequence of classical neural network layers (such as fully connected layers, convolutional layers, activation functions, batch normalization). The state trajectories of (1) are defined based on the solution $\Phi:\mathbb{R}\times\mathbb{R}^{n}\rightarrow\mathbb{R}^{n}$ of the corresponding initial value problem:
@@ 3.1 Lagrange remainder
Proposition 1(Lagrange remainder [25])
@@ 3.2 Error function
In our context of Section 2, the function $f$ is assumed not to be explicitly dependent on the depth $t$ due to its definition as a single residual block with classical layers. Therefore, the partial derivative $\frac{\partial f(x(t))}{\partial t}$ is equal to 0, and the third term of (5) thus reduces to:
@@ 3.3 Bounding the error set
$\mathcal{R}_{\varepsilon}(\mathcal{X}_{in})$ $=\left\{\Phi(1,u)-(u+f(u))~|~u\in\mathcal{X}_{in}\right\}$
$=\left\{\left.\frac{1}{2}f^{\prime}(\Phi(t^{*},u))f(\Phi(t^{*},u))~\right|~t^{*}\in[0,1],~u\in\mathcal{X}_{in}\right\}.$ (9)
@@
The error set can then be re-written based on the above reachable tube definition, by replacing $\Phi(t^{*},u)$ (with $t^{*}\in[0,1]$ and $u\in\mathcal{X}_{in}$) in (9) by $x\in\mathcal{R}^{\text{tube}}_{\text{neural ODE}}(\mathcal{X}_{in})$.
$\mathcal{R}_{\varepsilon}(\mathcal{X}_{in})$ $=\left\{\left.\frac{1}{2}f^{\prime}(x)f(x)~\right|~x\in\mathcal{R}^{\text{tube}}_{\text{neural ODE}}(\mathcal{X}_{in})\right\}$
@@
The next step, in line 2 of Algorithm 1, is to over-approximate this error set $\mathcal{R}_{\varepsilon}(\mathcal{X}_{in})$. One possible approach to achieve this is to define the static function $\varepsilon=\frac{1}{2}f^{\prime}(x)f(x)$ and apply to it some set-propagation techniques (such as interval arithmetic [12], Taylor models [19], or affine arithmetic [6]) to bound the set of output errors $\varepsilon$ corresponding to any state $x\in\Omega^{\text{tube}}_{\text{neural ODE}}(\mathcal{X}_{in})$ in the reachable tube over-approximation. An alternative approach, which provided a tighter error bounding set in the particular case of the numerical example presented in Section 4, is to define the discrete-time nonlinear system $x^{+}=\frac{1}{2}f^{\prime}(x)f(x)$, and then use existing reachability analysis toolboxes to over-approximate the reachable set of this system after one time step, which corresponds to bounding the image of the error function. Note that in this case, it is important that this final reachable set is computed as a single step, and not decomposed into a sequence of smaller intermediate steps whose iterative updates of the internal state would have no mathematical meaning for the static (stateless) error function.
@@
where $L$ is a Lipschitz constant of the neural ODE vector field. The term $\left\|\frac{1}{2}f^{\prime}(x)f(x)\right\|_{\infty}$ can be obtained by first over-approximating the error set by $\Omega_{\varepsilon}(\mathcal{X}_{in})$ in the same way we did, but the infinity norm forces to expand this set to make it symmetrical around 0, and then keeping only the maximum value among its components (thus corresponding to a second expansion of this set into an hypercube whose width along all dimensions is the largest width of the previous set). In addition, for any system with non-zero Lipschitz constant, the factor $\frac{e^{L}-1}{L}$ is always greater than $1$, which increases this error bound even more.
@@ 3.4 Verification proxy
We first focus on the case of Algorithm 1 to verify the safety property on the neural ODE, based on the reachability analysis of the ResNet. This first verification proxy relies on the set-based version of (6) using the Minkowski sum:
@@
stating that the reachable output set of the neural ODE is contained in the output set over-approximation of the ResNet $\Omega_{\text{ResNet}}(\mathcal{X}_{in})$, expanded by the bounding set of the error $\Omega_{\varepsilon}(\mathcal{X}_{in})$ obtained after applying the first two lines of Algorithm 1 as described in Section 3.3.
Therefore, this verification procedure is achieved as in Algorithm 1, by first using existing set-propagation or reachability analysis tools to compute an over-approximation $\color[rgb]{0,0,1}\Omega_{\text{ResNet}}(\mathcal{X}_{in})$} of the ResNet output set (line 3). Then in line 4, an over-approximation of the neural ODE output set can be deduced from (11) by taking the Minkowski sum of $\color[rgb]{0,0,1}\Omega_{\text{ResNet}}(\mathcal{X}_{in})$} and our error bound $\color[rgb]{1,0,0}\Omega_{\varepsilon}(\mathcal{X}_{in})$}. If $\Omega_{\text{neural ODE}}(\mathcal{X}_{in})$ is contained in the safe set $\color[rgb]{0,0,0}\mathcal{X}_{s}$}, then the neural ODE satisfies the safety property, otherwise the result is inconclusive (line 5-9).
@@
2: compute the over-approximation of the error set $\color[rgb]{1,0,0}\Omega_{\varepsilon}(\mathcal{X}_{in})$}, $\forall x\in\Omega^{\text{tube}}_{\text{neural ODE}}(\mathcal{X}_{in})$;
3: compute the over-approximation of the ResNet output $\color[rgb]{0,0,1}\Omega_{\text{ResNet}}(\mathcal{X}_{in})$};
@@
5:if $\Omega_{\text{neural ODE}}(\mathcal{X}_{in})\subseteq{\color[rgb]{0,0,0}\mathcal{X}_{s}}$ then
@@
2: compute the over-approximation of the negative error set $\color[rgb]{1,0,0}\Omega_{-\varepsilon}(\mathcal{X}_{in})$}, $\forall x\in\Omega^{\text{tube}}_{\text{neural ODE}}(\mathcal{X}_{in})$;
@@
5:if ${\color[rgb]{0,0,1}\Omega_{\text{ResNet}}(\mathcal{X}_{in})}\subseteq{\color[rgb]{0,0,0}\mathcal{X}_{s}}$ then
@@
The soundness of the verification framework is guaranteed because both algorithms rely on over-approximations of the true reachable sets. Specifically, (11) ensures that $\mathcal{R}_{\text{neural ODE}}(\mathcal{X}_{in})\subseteq\Omega_{\text{neural ODE}}(\mathcal{X}_{in})$, and (12) ensures ${\color[rgb]{0,0,1}\mathcal{R}_{\text{ResNet}}(\mathcal{X}_{in})}\subseteq{\color[rgb]{0,0,1}\Omega_{\text{ResNet}}(\mathcal{X}_{in})}$. These inclusions hold due to the conservative nature of the considered reachability analysis and error bound computations in Section 3.3 (Theorem 3.1).
@@ 4.2 Computing the error bound
Using CORA [1], we compute the error bound $\Omega_{\varepsilon}(\mathcal{X}_{in})$ from Theorem 3.1 as follows. First, we over-approximate the reachable tube of the neural ODE $\mathcal{R}^{\text{tube}}_{\text{neural ODE}}$ over the time interval $[0,1]$ as a sequence of zonotopes, where each zonotope corresponds to an intermediate time range. For each zonotope in the reachable tube, we bound the image of the error function (7) by applying a discrete-time reachability analysis method at $t=1$. This results in a new zonotope that over-approximates the error set starting from that particular reachable tube zonotope. The total error set is thus guaranteed to be contained in the union of these error zonotopes across all time steps. To simplify its use in the safety verification experiments in Section 4.3, we compute the interval hull of this union, yielding a hyperrectangle that over-approximates $\color[rgb]{1,0,0}\Omega_{\varepsilon}(\mathcal{X}_{in})$} illustrated in Figure 2 in red, and showing 20 error zonotopes in different colors, corresponding to the error bound of each intermediate time range used in the reachable tube.
@@
To contextualize our proposed error bound, we compare it with the error bound proposed in [26]. For that, we first compute the infinity norm of our error set $\|{\color[rgb]{1,0,0}\Omega_{\varepsilon}(\mathcal{X}_{in})}\|_{\infty}=0.064$, which corresponds to a positive and scalar bound on the error, thus implying that its set representation in the state space (represented in yellow in Figure 3) is necessarily symmetrical around 0 and with the width that is identical on all dimensions (since the infinity norm takes the largest width across all dimensions). The set-based error bound ($\color[rgb]{1,0,0}\Omega_{\varepsilon}(\mathcal{X}_{in})$} represented in red) obtained from our method is thus always contained in this infinity norm.
Next, we compute the Lipschitz constant for the vector field of the neural ODE $L=\|\tau+W\|_{\infty}=3.62$, and then we obtain the error bound in [26] as $\frac{(e^{L}-1)}{L}\|\Omega_{\varepsilon}(\mathcal{X}_{in})\|_{\infty}=0.64$. This final error bound, represented in magenta in Figure 3, is 10 times wider (on each dimension) than the infinity norm of our error set in yellow, and about $16$ millions times larger (in volume over the $5$-dimensional state space) than our error set $\color[rgb]{1,0,0}\Omega_{\varepsilon}(\mathcal{X}_{in})$} in red. The improved tightness of our proposed approach is therefore very significant.
@@ 4.3 Experiments on safety verification
Using the error bound computed in Section 4.2, we can verify safety properties for the neural ODE output set based on the ResNet output set and the error bound set (i.e., $\color[rgb]{1,0,0}\Omega_{\text{ResNet}}(\mathcal{X}_{in})+\Omega_{\varepsilon}(\mathcal{X}_{in})$}), or vice versa for the ResNet output set based on the neural ODE output set and the negative error bound set (i.e., $\color[rgb]{1,0,0}\Omega_{\text{neural ODE}}(\mathcal{X}_{in})+\Omega_{-\varepsilon}(\mathcal{X}_{in})$}).
In Figure 4, we compute the over-approximation of the ResNet output set $\color[rgb]{0,0,1}\Omega_{\text{ResNet}}$} using simple bound propagation through the ResNet function with CORA. By adding the error bound $\color[rgb]{1,0,0}\Omega_{\varepsilon}$}, we obtain a zonotope (shown in red) that is guaranteed to contain $\mathcal{R}_{\text{neural ODE}}(\mathcal{X}_{in})$. The figure also includes black points representing neural ODE outputs for random initial conditions in $\mathcal{X}_{in}$, with their convex hull (black set) approximating the true reachable set $\mathcal{R}_{\text{neural ODE}}(\mathcal{X}_{in})$. Since the safe set $\color[rgb]{0,0,0}\mathcal{X}_{s}$} contains the over-approximation $\color[rgb]{1,0,0}\Omega_{\text{ResNet}}(\mathcal{X}_{in})+\Omega_{\varepsilon}(\mathcal{X}_{in})$}, we guarantee that the neural ODE true reachable set is safe, as:
${\color[rgb]{0,0,0}\mathcal{X}_{s}}\supseteq{\color[rgb]{1,0,0}\Omega_{\text{ResNet}}(\mathcal{X}_{in})+\Omega_{\varepsilon}(\mathcal{X}_{in})}\supseteq\mathcal{R}_{\text{neural ODE}}.$
From Figure 4, we can see that the ResNet and neural ODE reachable sets are very similar due to the ResNet role as a discretization of the neural ODE, but they are not identical. Indeed, some neural ODE outputs (black points) lie outside $\color[rgb]{0,0,1}\Omega_{\text{ResNet}}$}, highlighting the necessity of the error bound $\color[rgb]{1,0,0}\Omega_{\varepsilon}(\mathcal{X}_{in})$} to ensure that the over-approximation captures all possible neural ODE outputs.
@@
Conversely, in Figure 5, we compute the over-approximation of the neural ODE reachable set $\color[rgb]{0,0,0}\Omega_{\text{neural ODE}}(\mathcal{X}_{in})$}. By adding the negative error bound $\color[rgb]{1,0,0}\Omega_{-\varepsilon}$}, we obtain a zonotope (shown in red) that encapsulates $\color[rgb]{0,0,1}\mathcal{R}_{\text{ResNet}}(\mathcal{X}_{in})$}. Similarly, the figure includes blue points representing ResNet outputs for random inputs in $\mathcal{X}_{in}$, with their convex hull (blue set) approximating the true reachable set $\color[rgb]{0,0,1}\mathcal{R}_{\text{ResNet}}(\mathcal{X}_{in})$}. Since the safe set $\color[rgb]{0,0,0}\mathcal{X}_{s}$} is a super set that contains the over-approximation $\color[rgb]{1,0,0}\Omega_{\text{neural ODE}}(\mathcal{X}_{in})+\Omega_{-\varepsilon}(\mathcal{X}_{in})$}, we guarantee that the ResNet true reachable set is safe, as:
${\color[rgb]{0,0,0}\mathcal{X}_{s}}\supseteq{\color[rgb]{1,0,0}\Omega_{\text{neural ODE}}(\mathcal{X}_{in})+\Omega_{-\varepsilon}(\mathcal{X}_{in})}\supseteq{\color[rgb]{0,0,1}\mathcal{R}_{\text{ResNet}}}.$